# itemsep=1pt,topsep=2pt,parsep=1pt,partopsep=1pt

### v2

@@ 1.4. Contributions and discussions
An effectively studied marginally relevant disordered model at criticality is the directed polymer as presented in the seminal work [CSZ23]. It proves that for the two-dimensional directed polymer with temperature in the critical window, the family of polymer measures (similar to (1.26)), indexed by starting and end times, converges in finite dimensional distributions to the critical 2d stochastic heat flow. The same convergence is also believed to hold for the critical 2d SHE, the continuous analogue of the directed polymer, although [CSZ23] does not provide proof details. This outcome offers an interpretation of the solution to a critical singular SPDE for the first time. The proof therein is robust enough, enabling us to adapt it for Theorem 1.7. Using the same framework, we provide a complete proof of Theorem 1.10 as well, which, in turn, complements the case of SHE. Besides the technical challenges needed to adapt the mathematical framework of [CSZ23] to the current setting, the interest of our paper lies in the establish of CDPM for the disordered pinning model, and in the applications to key problems in SPDEs and mathematical finance, which we highlight as follows
@@
  • Firstly, the limiting CDPM of the marginally relevant pinning model at the critical temperature is identified for the first time. For the pinning model (1.2), when the disorder is relevant, a scaling limit of the pinning partition function was obtained in [CSZ16]; when the disorder is marginally relevant and the temperature is in the so-called subcritical regime, a scaling limit of the pinning partition function was established in [CSZ17]. However, to the best of our knowledge, for marginally relevant pinning model at criticality, how to obtain a scaling limit was unknown in both physics and mathematics literature.
@@
  • Secondly, we give, for the first time, a meaning to the solution of the multiplicative SVE (1.29) with critical index $H=0$, in contrast to the rough volatility modelled by additive SVE with $H=0$ as studied in [NR18]. By connecting with SHE, we define a solution to such singular models of SDEs, thereby extending their scope from the subcritical phase to the critical phase.
@@
  • Thirdly, in Theorem 1.6 we establish the equivalence of convergence between random measures on time coordinate and on spatial coordinate, which allows us to use the techniques developed in [CSZ23] to deduce Theorem 1.1 by proving Theorem 1.7. In particular, as in (2.9)-(2.10), we show that the measures on time coordinate and on spatial coordinate can be expressed as weighted averages of each other, where the weights are probability densities for the first hitting times of a Brownian motion, which might be of independent interests. Although the equivalence seems intuitively natural, it turns out that the proof is highly non-trivial, crucially relying on the expressions of transition kernel related to the Brownian motion, see, e.g., (3.15).
@@
  • Fourthly, the paper can be read as a complement to [CSZ23]. Some proofs that were sketched in [CSZ23] are now provided in full detail, such as the moment computations of the coarse-grained model in Section 7. We also provide enough details in proving Theorem 1.10, the case of SHE, demonstrating the robustness of techniques in [CSZ23] under continuous settings.
@@
In the next section, we will illustrate our proof strategy for Theorem 1.6. For Theorems 1.7 and 1.10, we will first review the strategies in [CSZ23], and then discuss the adaption to our case. Along the way, we will also introduce some preliminary ingredients that will be moreover used throughout the paper. At the end of that section, we will state the organization of the rest of the paper.
2. Proof Outlines, Notations and Tools
@@
Note that Theorem 1.1 is a direct consequence of Theorem 1.6 and Theorem 1.7. Theorem 1.7 and Theorem 1.10 are proved in the same framework as developed in [CSZ23]. We thereby only focus on explaining the proofs for Theorem 1.6 and Theorem 1.7. Before illustrating the proof for Theorem 1.7, we will first review the framework in [CSZ23], so that it is more straightforward to explain the modifications needed when adapting. We will also introduce some necessary notations and tools, including polynomial chaos expansion, the Dickman subordinator, coarse-grained disorders and a coarse-grained model.
2.1. Proof outline for Theorem 1.6
@@
We need to show that the random pinning measure ${\mathcal{Z}}_{N}^{\beta_{N}}({\mathrm{d}}s,{\mathrm{d}}t)$ (resp. the random polymer measure $\widetilde{\mathcal{Z}}_{0,N}^{\beta_{N}}({\mathrm{d}}x,{\mathrm{d}}y)$) has a unique limit $\mathscr{L}^{\vartheta}({\mathrm{d}}s,{\mathrm{d}}t)$ (resp. ~$\widetilde{\mathscr{L}}^{\vartheta}_{0,1}({\mathrm{d}}x,{\mathrm{d}}y)$). Thanks to [Kal17, Corollary 4.14], it suffices to apply test functions $h\in C_{c}([0,1]^{2}_{\leq})$ (resp. $\phi\in C_{c}({\mathbbm{R}}^{2})$) w.r.t. the random measures ${\mathcal{Z}}_{N}^{\beta_{N}}({\mathrm{d}}s,{\mathrm{d}}t)$ (resp. $\widetilde{{\mathcal{Z}}}_{0,N}^{\beta_{N}}({\mathrm{d}}x,{\mathrm{d}}y)$), and show that the resulting sequence of random variables has a weak limit. In particular, to prove Theorem 1.6, we only need to consider test functions of the form $f(s)g(t)$ with $f,g\in C_{c}([0,1])$ (resp. $\varphi(x)\psi(y)$ with $\varphi,\psi\in C_{c}({\mathbbm{R}})$) by a density argument, and show that the following two statements imply each other:
(2.1) ${\mathcal{Z}}_{N}^{\beta_{N}}(f,g)$ $:=\iint_{[0,1]^{2}_{\leq}}f(s){\mathcal{Z}}_{N}^{\beta_{N}}({\mathrm{d}}s,{\mathrm{d}}t)g(t)\quad\text{converges weakly for all}~{}f,g\in C_{c}([0,1]);$
@@
(2.2) $\widetilde{\mathcal{Z}}_{0,N}^{\beta_{N}}(\varphi,\psi)$ $:=\iint_{{\mathbbm{R}}^{2}}\varphi(x)\widetilde{\mathcal{Z}}_{0,N}^{\beta_{N}}({\mathrm{d}}x,{\mathrm{d}}y)\psi(y)\quad\text{converges weakly for all}~{}\varphi,\psi\in C_{c}({\mathbbm{R}}).$
A key step is to compare decompositions of ${\mathcal{Z}}_{N}^{\beta_{N}}(f,g)$ and $\widetilde{\mathcal{Z}}_{0,N}^{\beta_{N}}(\varphi,\psi)$.
@@
We first decompose $\widetilde{\mathcal{Z}}_{0,N}^{\beta_{N}}(\varphi,\psi)$ according to the first and last hitting times to 00 for the random walk in the time interval $[1,N-1]$. Let the first and last hitting times be
(2.3) $\tau_{1}$ $:=\inf\{n\geq 1:S_{n}=0~{}\text{and}~{}S_{k}\neq 0~{}\text{for}~{}k=1,\cdots,n-1\},$
(2.4) $\tau_{2}$ $:=\tau_{2}^{(N)}=\sup\{n\leq N-1:S_{n}=0~{}\text{and}~{}S_{k}\neq 0~{}\text{for}~{}k=n+1,\cdots,N-1\}.$
@@
(2.5) $\widetilde{Z}_{0,N}^{\omega,\beta_{N}}(x,y)=\widetilde{Z}_{0,N}^{\omega,\beta}(x,y)\left[\mathbbm{1}_{\{\tau_{1}\geq N\}}\right]+\sum\limits_{m=1}^{N-1}\sum\limits_{n=m}^{N-1}\widetilde{Z}_{0,N}^{\omega,\beta_{N}}(x,y)\left[\mathbbm{1}_{\{\tau_{1}=m,\tau_{2}=n\}}\right],$
@@
(2.6) $\widetilde{Z}_{0,N}^{\omega,\beta}(x,y)\left[\mathbbm{1}_{\{\tau_{1}\geq N\}}\right]={\mathbf{P}}^{x}(S_{N}=y,\tau_{1}\geq N),$
and due to the absence of the disorders at times 00 and $N$,
(2.7) $\widetilde{Z}_{0,N}^{\omega,\beta}(x,y)\left[\mathbbm{1}_{\{\tau_{1}=m,\tau_{2}=n\}}\right]={\mathbf{P}}^{x}(\tau_{1}=m)\times Z^{\omega,\beta_{N}}_{m,n}\times{\mathbf{P}}^{x}(\tau_{2}=n,S_{N}=y|S_{n}=0),$
@@
(2.8) $q_{x}(n):={\mathbf{P}}^{x}(\tau_{1}=n)$
to denote the probability of the first hitting time to 00 for a random walk starting at $x$. Note that by the Markov property and a time reversal, we have that
@@
(2.9) $\begin{split}&\widetilde{\mathcal{Z}}_{0,N}^{\beta_{N}}(\varphi,\psi)=\iint_{{\mathbbm{R}}\times{\mathbbm{R}}}\varphi(x)\sqrt{N}\hskip 1.0pt{\mathbf{P}}^{x_{N}}(S_{N}=y_{N},\tau_{1}\geq N)\psi(y){\mathrm{d}}x{\mathrm{d}}y\\ +\sum\limits_{m=1}^{N-1}\sum\limits_{n=m}^{N-1}&\Big{(}\int_{{\mathbbm{R}}}\varphi(x)\sqrt{N}q_{x_{N}}(m){\mathrm{d}}x\Big{)}\Big{(}\sqrt{N}Z^{\omega,\beta_{N}}_{m,n}\Big{)}\Big{(}\int_{{\mathbbm{R}}}\psi(x)\sqrt{N}q_{-y_{N}}(N-n){\mathrm{d}}y\Big{)},\end{split}$
@@
(2.10) ${\mathcal{Z}}_{N}^{\beta_{N}}(f,g)=\sum\limits_{m=0}^{N-1}\sum\limits_{n=m}^{N-1}\Big{(}\int_{\frac{m}{N}}^{\frac{m+1}{N}}f(s){\mathrm{d}}s\Big{)}\Big{(}\sqrt{N}Z_{m,n}^{\omega,\beta_{N}}\Big{)}\Big{(}\int_{\frac{n}{N}}^{\frac{n+1}{N}}g(t){\mathrm{d}}t\Big{)}.$
Note that the first term at the right-hand side of (2.9) is deterministic. To show (i)$\Longleftrightarrow$(ii) in Theorem 1.7, it suffices to show that (a) the deterministic term converges; (b) the term $m=0$ in (2.10) is negligible; (c) the two double-sums with $11\leq m\leq n\leq N-$ 1 in (2.10) and (2.9) can be made arbitrarily close by choosing suitable $f,g,\varphi,\psi$. While (a) and (b) are somewhat direct to check, (c) is verified through a crucial observation (3.15) regarding the heat kernel, that allows approximations between test functions $f,g$ and $\varphi,\psi$.
2.2. Proof strategy in [CSZ23]
In [CSZ23], the authors treated the classic directed polymer model, whose point-to-point partition function is defined by
@@ 2.2. Proof strategy in [CSZ23]
where the disorders $\omega_{n,x}$ are located on the whole time-space lattice ×${\mathbbm{N}}\times{\mathbbm{Z}}^{2}$. The related random measure is defined similarly to (1.26), by
@@
(2.11) ${\mathcal{B}}_{N,\varepsilon}(\mathrm{i},\mathrm{a}):=\big{(}(\mathrm{i}-1)\varepsilon N,\mathrm{i}\varepsilon N\big{]}\times\big{(}(\mathrm{a}-(1,1))\sqrt{\varepsilon N},\mathrm{a}\sqrt{\varepsilon N}\big{]}:={\mathcal{T}}_{N,\varepsilon}(\mathrm{i})\times{\mathcal{S}}_{N,\varepsilon}(\mathrm{a}),\quad\mathrm{i}\in{\mathbbm{N}},\mathrm{a}\in{\mathbbm{Z}}^{2},$
@@
Finally, to show that ${\mathbbm{E}}[f(\overline{\mathcal{Z}}_{N})]$ is a Cauchy sequence, we can first replace $\overline{\mathcal{Z}}_{N}$ there by the coarse-grained model to obtain a new expectation sequence, and then show that the new expectation sequence is a Cauchy sequence as $\varepsilon\to 0$, thanks to an enhanced Lindeberg principle (see Appendix A). It is worth mentioning that in this last step, we need to estimate the higher moments of the coarse-grained disorders. This is done by estimating the higher moments of $\overline{\mathcal{Z}}_{N}$, which is a byproduct of [CSZ23].
2.3. Proof outlines for Theorem 1.7 and Theorem 1.10
@@
  • In our variant of directed polymer model (1.25), the disorders are only located on the time-space line ×${\mathbbm{Z}}\times\{0\}$. Hence, except when visiting the origin, the random walk distribution is not tilted by the polymer measure. Therefore, we only need to introduce the mesoscopic time interval ${\mathcal{T}}_{N,\varepsilon}(\mathrm{i})$ in (2.11), instead of the mesoscopic time-space boxes ${\mathcal{B}}_{N,\varepsilon}(\mathrm{i},\mathrm{a})$, since we only concern the return times to 00.
  • As a consequence, we have much less estimates on random walk transition probability than those in [CSZ23]. To be specific, unlike in [CSZ23], large deviation estimates are not necessary in our case. We only require a local limit theorem, and in particular, a sharp estimate for the probability of the random walk returning to 0. This allows us to relax the constraint on the random walk to the assumption (1.10) (e.g. see (4.27)-(4.28)).
To prove Theorem 1.10, we adapt the techniques for continuous setting in [CSZ19b] and show that $\iint_{{\mathbbm{R}}^{2}}\varphi(x)u_{0,1}^{\delta}({\mathrm{d}}x,{\mathrm{d}}y)\psi(y)$ (see Remark 1.11) can also be approximated by the coarse-grained model (see 2.46) introduced when proving Theorem 1.7.
2.4. Polynomial chaos expansion
The polynomial chaos expansion is a celebrated tool in treating random polymer models when the random environment is independent in time, see, e.g., [CSZ16, CSZ17, CSZ23]. We illustrate it by taking $Z_{N,\beta_{N}}^{\omega}$ defined in (1.13) as an example.

### v3

@@ 1.4 Contributions and discussions
An effectively studied marginally relevant disordered model at criticality is the directed polymer as presented in the seminal work [15]. It proves that for the two-dimensional directed polymer with temperature in the critical window, the family of polymer measures (similar to (1.26)), indexed by starting and end times, converges in finite dimensional distributions to the critical 2d stochastic heat flow. The same convergence is also believed to hold for the critical 2d SHE, the continuous analogue of the directed polymer, although [15] does not provide proof details. This outcome offers an interpretation of the solution to a critical singular SPDE for the first time. The proof therein is robust enough, enabling us to adapt it for Theorem 1.7. Using the same framework, we provide a complete proof of Theorem 1.10 as well, which, in turn, complements the case of SHE. Besides the technical challenges needed to adapt the mathematical framework of [15] to the current setting, the interest of our paper lies in the establish of the CDPM for the disordered pinning model, and in the applications to key problems in SPDEs and mathematical finance, which we highlight as follows:
@@
  • Firstly, the limiting CDPM of the marginally relevant pinning model at the critical temperature is identified for the first time. For the pinning model (1.2), when the disorder is relevant, a scaling limit of the pinning partition function was obtained in [11]; when the disorder is marginally relevant and the temperature is in the so-called subcritical regime, a scaling limit of the pinning partition function was established in [12]. However, to the best of our knowledge, for marginally relevant pinning model at criticality, how to obtain a scaling limit was unknown in both physics and mathematics literature.
@@
  • Secondly, we give, for the first time, a meaning to the solution of the multiplicative SVE (1.29) with critical index $H=0$, in contrast to the rough volatility modelled by additive SVE with $H=0$ as studied in [32]. By connecting with SHE, we define a solution to such singular models of SDEs, thereby extending their scope from the subcritical phase to the critical phase.
@@
  • Thirdly, in Theorem 1.6 we establish the equivalence of convergence between random measures on time coordinate and on spatial coordinate, which allows us to use the techniques developed in [15] to deduce Theorem 1.1 by proving Theorem 1.7. In particular, as in (2.9)-(2.10), we show that the measures on time coordinate and on spatial coordinate can be expressed as weighted averages of each other, where the weights are probability densities for the first hitting times of a Brownian motion, which might be of independent interests. Although the equivalence seems intuitively natural, it turns out that the proof is highly non-trivial, crucially relying on the expressions of transition kernel related to the Brownian motion, see, e.g., (3.15).
@@
  • Fourthly, the paper can be read as a complement to [15]. Some proofs that were sketched in [15] are now provided in more details, such as the moment computations of the coarse-grained model w.r.t. disorders $W_{s,t,u}^{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}}$ in Section 7 (see Lemma 7.1). We also provide enough details in proving Theorem 1.10, the SHE case, demonstrating the robustness of techniques in [15] under continuous settings.
@@
In the next section, we will illustrate our proof strategy for Theorem 1.6. For Theorems 1.7 and 1.10, we will first review the strategies in [15], and then discuss the adaption to our case. Along the way, we will also introduce some preliminary ingredients that will be moreover used throughout the paper. At the end of that section, we will state the organization of the rest of the paper.
2 Proof Outlines, Notations and Tools
@@
Note that Theorem 1.1 is a direct consequence of Theorem 1.6 and Theorem 1.7. Theorem 1.7 and Theorem 1.10 are proved in the same framework as developed in [15]. We thereby only focus on explaining the proofs for Theorem 1.6 and Theorem 1.7. Before illustrating the proof of Theorem 1.7, we will first review the framework in [15], so that it is more straightforward to explain the modifications needed when adapting. We will also introduce some necessary notations and tools, including polynomial chaos expansion, the Dickman subordinator, coarse-grained disorders and a coarse-grained model.
2.1 Proof outline for Theorem 1.6
@@
We need to show that the random pinning measure ${\mathcal{Z}}_{N}^{\beta_{N}}({\mathrm{d}}s,{\mathrm{d}}t)$ (resp. the random polymer measure $\widetilde{\mathcal{Z}}_{0,N}^{\beta_{N}}({\mathrm{d}}x,{\mathrm{d}}y)$) has a unique limit $\mathscr{L}^{\vartheta}({\mathrm{d}}s,{\mathrm{d}}t)$ (resp. $\widetilde{\mathscr{L}}^{\vartheta}_{0,1}({\mathrm{d}}x,{\mathrm{d}}y)$). Thanks to [28, Corollary 4.14], it suffices to apply test functions $h\in C_{c}([0,1]^{2}_{\leq})$ (resp. $\phi\in C_{c}({\mathbbm{R}}^{2})$) w.r.t. the random measures ${\mathcal{Z}}_{N}^{\beta_{N}}({\mathrm{d}}s,{\mathrm{d}}t)$ (resp. $\widetilde{{\mathcal{Z}}}_{0,N}^{\beta_{N}}({\mathrm{d}}x,{\mathrm{d}}y)$), and show that the resulting sequence of random variables has a weak limit. In particular, to prove Theorem 1.6, we only need to consider test functions of the form $f(s)g(t)$ with $f,g\in C_{c}([0,1])$ (resp. $\varphi(x)\psi(y)$ with $\varphi,\psi\in C_{c}({\mathbbm{R}})$) by a density argument, and show that the following two statements imply each other:
${\mathcal{Z}}_{N}^{\beta_{N}}(f,g):=\iint_{[0,1]^{2}_{\leq}}f(s){\mathcal{Z}}_{N}^{\beta_{N}}({\mathrm{d}}s,{\mathrm{d}}t)g(t)~~\text{converges weakly for all}~f,g\in C_{c}([0,1]);$ (2.1)
@@
$\widetilde{\mathcal{Z}}_{0,N}^{\beta_{N}}(\varphi,\psi):=\iint_{{\mathbbm{R}}^{2}}\varphi(x)\widetilde{\mathcal{Z}}_{0,N}^{\beta_{N}}({\mathrm{d}}x,{\mathrm{d}}y)\psi(y)~~\text{converges weakly for all}~\varphi,\psi\in C_{c}({\mathbbm{R}}).$ (2.2)
A key step is to compare the decompositions of ${\mathcal{Z}}_{N}^{\beta_{N}}(f,g)$ and $\widetilde{\mathcal{Z}}_{0,N}^{\beta_{N}}(\varphi,\psi)$.
@@
We first decompose $\widetilde{\mathcal{Z}}_{0,N}^{\beta_{N}}(\varphi,\psi)$ according to the first and last hitting times to 0 for the random walk in the time interval [1,N1][1,N-1]. Let the first and last hitting times be
$\tau_{1}$ $:=\inf\{n\geq 1:S_{n}=0~\text{and}~S_{k}\neq 0~\text{for}~k=1,\cdots,n-1\},$ (2.3)
$\tau_{2}$ $:=\tau_{2}^{(N)}=\sup\{n\leq N-1:S_{n}=0~\text{and}~S_{k}\neq 0~\text{for}~k=n+1,\cdots,N-1\}.$ (2.4)
@@
$\widetilde{Z}_{0,N}^{\omega,\beta_{N}}(x,y)=\widetilde{Z}_{0,N}^{\omega,\beta}(x,y)\left[\mathbbm{1}_{\{\tau_{1}\geq N\}}\right]+\sum\limits_{m=1}^{N-1}\sum\limits_{n=m}^{N-1}\widetilde{Z}_{0,N}^{\omega,\beta_{N}}(x,y)\left[\mathbbm{1}_{\{\tau_{1}=m,\tau_{2}=n\}}\right],$ (2.5)
@@
$\widetilde{Z}_{0,N}^{\omega,\beta}(x,y)\left[\mathbbm{1}_{\{\tau_{1}\geq N\}}\right]={\mathbf{P}}^{x}(S_{N}=y,\tau_{1}\geq N),$ (2.6)
and due to the absence of the disorders at times 0 and $N$,
$\widetilde{Z}_{0,N}^{\omega,\beta}(x,y)\left[\mathbbm{1}_{\{\tau_{1}=m,\tau_{2}=n\}}\right]={\mathbf{P}}^{x}(\tau_{1}=m)\times Z^{\omega,\beta_{N}}_{m,n}\times{\mathbf{P}}^{x}(\tau_{2}=n,S_{N}=y|S_{n}=0),$ (2.7)
@@
$q_{x}(n):={\mathbf{P}}^{x}(\tau_{1}=n)$ (2.8)
to denote the probability of the first hitting time to 0 for a random walk starting at $x$. Note that by the Markov property and a time reversal, we have that
@@
$\begin{split}&\widetilde{\mathcal{Z}}_{0,N}^{\beta_{N}}(\varphi,\psi)=\iint_{{\mathbbm{R}}\times{\mathbbm{R}}}\varphi(x)\sqrt{N}\hskip 1.0pt{\mathbf{P}}^{x_{N}}(S_{N}=y_{N},\tau_{1}\geq N)\psi(y){\mathrm{d}}x{\mathrm{d}}y\\ +\sum\limits_{m=1}^{N-1}\sum\limits_{n=m}^{N-1}&\Big(\int_{{\mathbbm{R}}}\varphi(x)\sqrt{N}q_{x_{N}}(m){\mathrm{d}}x\Big)\Big(\sqrt{N}Z^{\omega,\beta_{N}}_{m,n}\Big)\Big(\int_{{\mathbbm{R}}}\psi(x)\sqrt{N}q_{-y_{N}}(N-n){\mathrm{d}}y\Big),\end{split}$ (2.9)
@@
${\mathcal{Z}}_{N}^{\beta_{N}}(f,g)=\sum\limits_{m=0}^{N-1}\sum\limits_{n=m}^{N-1}\Big(\int_{\frac{m}{N}}^{\frac{m+1}{N}}f(s){\mathrm{d}}s\Big)\Big(\sqrt{N}Z_{m,n}^{\omega,\beta_{N}}\Big)\Big(\int_{\frac{n}{N}}^{\frac{n+1}{N}}g(t){\mathrm{d}}t\Big).$ (2.10)
[…]
2.2 Proof strategy in [15]
In [15], the authors treated the classic directed polymer model, whose point-to-point partition function is defined by
@@ 2.2 Proof strategy in [15]
where the disorders $\omega_{n,x}$ are located on the whole time-space lattice ${\mathbbm{N}}\times{\mathbbm{Z}}^{2}$. The related random measure is defined similarly to (1.26), by
@@
$\begin{split}{\mathcal{B}}_{N,\varepsilon}(\mathrm{i},\mathrm{a})&:=\big((\mathrm{i}-1)\varepsilon N,\mathrm{i}\varepsilon N\big]\times\big((\mathrm{a}-(1,1))\sqrt{\varepsilon N},\mathrm{a}\sqrt{\varepsilon N}\big]\\ &:={\mathcal{T}}_{N,\varepsilon}(\mathrm{i})\times{\mathcal{S}}_{N,\varepsilon}(\mathrm{a}),\quad\mathrm{i}\in{\mathbbm{N}},\mathrm{a}\in{\mathbbm{Z}}^{2},\end{split}$ (2.11)
@@
Finally, to show that ${\mathbbm{E}}[f(\overline{{\mathcal{Z}}}_{N})]$ is a Cauchy sequence, we can first replace $\overline{{\mathcal{Z}}}_{N}$ there by the coarse-grained model to obtain a new expectation sequence, and then show that the new expectation sequence is a Cauchy sequence as $\varepsilon\to 0$, thanks to an enhanced Lindeberg principle (see Appendix A). It is worth mentioning that in this last step, we need to estimate the higher moments of the coarse-grained disorders. This is done by estimating the higher moments of $\overline{{\mathcal{Z}}}_{N}$, which is a byproduct of [15].
2.3 Proof outlines for Theorem 1.7 and Theorem 1.10
@@
  • In our variant of directed polymer model (1.25), the disorders are only located on the time-space line ${\mathbbm{Z}}\times\{0\}$. Hence, except when visiting the origin, the random walk distribution is not tilted by the polymer measure. Therefore, we only need to introduce the mesoscopic time interval ${\mathcal{T}}_{N,\varepsilon}(\mathrm{i})$ in (2.11), instead of the mesoscopic time-space boxes ${\mathcal{B}}_{N,\varepsilon}(\mathrm{i},\mathrm{a})$, since we only concern the return times to 0.
  • As a consequence, we need less estimates on random walk transition probability than those in [15]. To be specific, unlike in [15], large deviation estimates are not necessary in our case. We only require a local limit theorem whose error satisfies some summable conditions. This allows us to relax the constraint on the random walk to the assumption (1.10) (e.g. see (4.27)-(4.28) and (5.13)).
To prove Theorem 1.10, we adapt the techniques for continuous setting in [14] and show that $\iint_{{\mathbbm{R}}^{2}}\varphi(x)u_{0,1}^{\delta}({\mathrm{d}}x,{\mathrm{d}}y)\psi(y)$ (see Remark 1.11) can also be approximated by the coarse-grained model (see 2.46) introduced when proving Theorem 1.7.
2.4 Polynomial chaos expansion
The polynomial chaos expansion is a celebrated tool in treating random polymer models when the random environment is independent in time, see, e.g., [11, 12, 15]. We illustrate it by taking $Z_{N,\beta_{N}}^{\omega}$ defined in (1.13) as an example.